OpenStudy (anonymous):

gallium has two isotopes. one isotope is 69-relative atomic mass and atomic number-31 Ga and has a relative abundance of 60.50%. The relative atomic mass of Gallium is 69.70. Find the relative isotopic mass of the other isotopes.

OpenStudy (abb0t):

well, the other one takes up 100 -60.5 = 39.5% so, \[(69.31 \times \frac{ 60.5 }{ 100 })+(y \times \frac{ 39 }{ 100 }) = 69.70\] y = answer

OpenStudy (anonymous):

but the answer in my textbook is 70.77 and i get y=2.55

OpenStudy (abb0t):

When I plugged in everything, I got that answer (rounded up). I used a very old calculator. But essentially, the same number.

OpenStudy (abb0t):

Make sure you are familiar with your calculator(s). They are powerful tools that know their stuff, it's up to you to know how to use it :)